# The Impact of Feature Scaling In Machine Learning: Effects on Regression and Classification Tasks

### v1

@@ III-C3 Z-score Normalization (ZSN)
where $X$ is the original feature value, $\mu$ is the mean of the feature, $\sigma$ is the standard deviation [^1] For transparency, we clarify that the population standard deviation was used throughout this study. of the feature, and $X_{\text{norm}}$ is the scaled value.
@@ IV Experiments
This section presents the empirical results of our extensive experiments. We selected five models and three scaling techniques, along with a baseline without scaling, as this subset already allows us to draw meaningful conclusions and observe differences between models. The complete results are available in the Appendix A.
@@ IV-A Impact on Validation Metrics
As anticipated, a key finding from our classification experiments is the differential impact of feature scaling on model performance. Table III shows the accuracy results. Ensemble methods, including Random Forest and the gradient boosting family (LightGBM, CatBoost, XGBoost), proved robust, delivering consistently high validation metrics irrespective of the preprocessing strategy. This inherent robustness offers a significant practical advantage, particularly in resource-constrained environments, since omitting the scaling step eliminates the associated memory and computational overhead. The Naive Bayes model showed similar scaling resistance, although its overall accuracy was not competitive with these top-tier ensembles. In stark contrast, the performance of Logistic Regression (LR), Support Vector Machines (SVM), K-Nearest Neighbor (KNN), TabNet and Multi-Layer Perceptrons (MLP) was highly dependent on the choice of scaler, revealing their pronounced sensitivity to data preprocessing through significant fluctuations in performance.
@@
IV-B Train and Inference time Impact
@@
The application of different scaling techniques had a variable impact on inference times in all models. For training time, the effects of scaling largely mirrored those seen in the validation accuracy results, as shown in Tables VII and IX. Certain models, notably tree-based ensembles, did not derive a significant speed benefit from feature scaling, while others were more sensitive. The preprocessing step for inference generally introduced a small, non-uniform computational cost, which could become more pronounced for large datasets or in applications requiring real-time inference.
@@ IV-C Results of Memory Usage (kB)
The analysis showed a clear distinction in memory consumption between scaling techniques. As expected, applying no scaling resulted in zero additional memory consumption, only the consumption to load the data. Among the actual scaling methods, the RobustScaler, StandardScaler,Tanh Transformer, and Hyperbolic Tangent were found to be the most memory intensive. In contrast, the MaxAbsScaler, MinMaxScaler, and Decimal Scaler consistently registered the lowest memory usage as shows in Table XI.
@@ V Limitation
  • Hyperparameter Optimization: The Machine Learning models in this study were evaluated using their default hyperparameters, as stated in the methodology. A comprehensive hyperparameter tuning process for each model-scaler-dataset combination was beyond the current scope but could potentially reveal different optimal pairings or further enhance model performance.
@@
  • Evaluation Metrics for Classification: The primary metric for classification tasks was accuracy. Although acknowledged as potentially misleading for imbalanced datasets, future work could incorporate a broader suite of metrics, such as F1 score, precision-recall AUC, or balanced precision, to provide a more nuanced understanding of performance, especially on datasets with skewed class distributions.
  • Dataset Size and Synthetic Data: For some of the smaller datasets utilized, the exploration of techniques like synthetic data generation or data augmentation was not performed. Such methods could potentially improve the robustness and performance of certain models, particularly data-hungry algorithms, and would be an interesting direction for further study.
@@ VI Conclusion
This comprehensive empirical study investigated the impact of 12 feature scaling techniques on 13 Machine Learning algorithms in 16 classification and regression datasets. Key findings reaffirmed the robustness of ensemble methods (e.g., Random Forest, gradient boosting family), which, along with Naive Bayes, largely maintained high performance irrespective of scaling. This offers efficiency gains by potentially avoiding preprocessing overhead. In stark contrast, models such as logistic regression, SVMs, MLPs, K-Nearest Neighbor and TabNet demonstrated high sensitivity, with their performance critically dependent on scaler choice - a pattern consistent across both task types. Computational analysis also indicated that scaling choices can influence training/inference times and memory usage, with certain scalers being notably more resource-intensive.
This study contributes with one of the first systematic evaluations of such an extensive array of models, some less common models like TabNet, and scaling techniques—including transformations, e.g Tanh (TT) and Hyperbolic Tangent (HT), which are less commonly benchmarked as general-purpose scalers, all within a unified Python framework. This is particularly relevant given that feature scaling is often applied in the literature without clear rationale, sometimes incorrectly before data splitting, which can lead to data leakage, or without verifying algorithm-specific benefits. By providing broad empirical evidence, our work offers clear guidance on how to mitigate these common issues, promoting informed scaling selection and more rigorous experimental design.

### v2

@@ III-C3 Z-score Normalization (ZSN)
where $X$ is the original feature value, $\mu$ is the mean of the feature, $\sigma$ is the standard deviation of the feature, and $X_{\text{norm}}$ is the scaled value.
@@ IV Experiments
This section presents the empirical results of our extensive experiments. We selected five representative models and three scaling techniques, along with a baseline without scaling, as this subset already allows us to draw meaningful conclusions and observe distinctions among models. The complete results are available in the Appendix A.
@@ IV-A Impact on Validation Metrics
As anticipated, one of the key findings from our classification experiments is the differential impact of feature scaling on model performance. Table III shows the accuracy results. Ensemble methods, including Random Forest and the gradient boosting family (LightGBM, CatBoost, XGBoost), demonstrated strong robustness by consistently achieving high validation performance irrespective of the preprocessing strategy or dataset. This inherent robustness offers a significant practical advantage, particularly in resource-constrained environments, since omitting the scaling step eliminates the associated memory and computational overhead. The Naive Bayes model showed similar scaling resistance, although its overall accuracy was not competitive with these top-tier ensembles. In stark contrast, the performance of Logistic Regression (LR), Support Vector Machines (SVM), K-Nearest Neighbor (KNN), TabNet, and Multi-Layer Perceptrons (MLP) was highly dependent on the choice of scaler, revealing their pronounced sensitivity to data preprocessing through significant fluctuations in performance.
@@
IV-B Impact on Training and Inference Times
@@
The application of different scaling techniques had a variable impact on inference times across the evaluated models. Notably, Classification and Regression Trees (CART) exhibited exceptionally robust behavior, remaining unaffected across all scaling methods and datasets. While certain Machine Learning algorithms — such as K-Nearest Neighbors (KNN), Random Forest, Support Vector Machine (SVM), and Support Vector Regressor (SVR) — showed more evident sensitivity to the choice of scaling technique, this was not the norm. For the majority of models, the preprocessing step introduced only a small, non-uniform computational overhead — as illustrated in Tables VIII and X — which may become more significant in the context of large-scale datasets or time-sensitive applications requiring real-time inference.
For training time, the effects of scaling largely mirrored those seen in the validation accuracy results, as shown in Tables VII and IX. Certain models, notably tree-based ensembles, did not derive a significant speed benefit from feature scaling, while others were more sensitive.
@@ IV-C Results of Memory Usage (kB)
The analysis revealed a clear distinction in memory consumption between scaling techniques. As expected, applying no scaling resulted in zero additional memory consumption, only the consumption to load the data. Among the actual scaling methods, the RobustScaler, StandardScaler, Tanh Transformer, and Hyperbolic Tangent were found to be the most memory-intensive. In contrast, the MaxAbsScaler, MinMaxScaler, and Decimal Scaler consistently registered the lowest memory usage as shown in Table XI.
@@ V Limitation
  • Hyperparameter Optimization: The Machine Learning models analyzed in this study were evaluated using their default hyperparameters, as outlined in the methodology. A comprehensive hyperparameter tuning process for each model–scaler–dataset combination was beyond the current scope; however, such optimization could potentially uncover different optimal pairings or further improve model performance.
@@
  • Evaluation Metrics for Classification: The primary metric for classification tasks was accuracy. Although acknowledged as potentially misleading for imbalanced datasets, future work could incorporate a broader suite of metrics, such as F1 score, precision, recall AUC, or balanced precision, to provide a more nuanced understanding of performance, especially on datasets with skewed class distributions.
  • Dataset Size and Synthetic Data: For some of the smaller datasets utilized, the exploration of techniques such as synthetic data generation or data augmentation was not performed. Such methods could potentially improve the robustness and performance of certain models, representing a promising direction for further research.
@@ VI Conclusion
This comprehensive empirical study investigated the impact of 12 feature scaling techniques on 14 Machine Learning algorithms in 16 classification and regression datasets. Key findings reaffirmed the robustness of ensemble methods (e.g., Random Forest, gradient boosting family), which, along with Naive Bayes, largely maintained high performance irrespective of scaling. This offers efficiency gains by potentially avoiding preprocessing overhead. In stark contrast, models such as Logistic Regression, SVMs, MLPs, K-Nearest Neighbor, and TabNet demonstrated high sensitivity, with their performance critically dependent on scaler choice — a pattern consistent across both task types. Computational analysis also indicated that scaling choices can influence training/inference times and memory usage, with certain scalers being notably more resource-intensive.
This study contributes with one of the first systematic evaluations of such an extensive array of models, some less common models like TabNet, and scaling techniques — including transformations, e.g Tanh (TT) and Hyperbolic Tangent (HT), which are less commonly benchmarked as general-purpose scalers —, all within a unified Python framework. This is particularly relevant given that feature scaling is often applied in the literature without clear rationale, sometimes incorrectly before data splitting, which can lead to data leakage, or without verifying algorithm-specific benefits. By providing broad empirical evidence, our work offers clear guidance on how to mitigate these common issues, promoting informed scaling selection and more rigorous experimental design.